### v1

@@ Abstract
Averaging dynamics drives countless processes in physics, biology, engineering, and the social sciences. In recent years, the $s$-energy has emerged as a useful tool for bounding the convergence rates of time-varying averaging systems. We derive new bounds on the $s$-energy, which we use to resolve a number of open questions in the areas of bird flocking, opinion dynamics, and distributed motion coordination. We also use our results to provide a theoretical validation for the idea of the “Overton Window” as an attracting manifold of viable group opinions. Our new bounds on the $s$-energy highlight its dependency on the connectivity of the underlying networks. In this vein, we use the $s$-energy to explain the exponential gap in the convergence rates of stationary and time-varying consensus systems.
@@ 1 Introduction
Dynamics based on local averaging has been the object of considerable attention [7, 9, 10, 11, 12, 13, 29, 34, 35, 36, 37, 39, 40, 43, 44, 45, 46, 49, 51, 52, 55, 62]. The reason is the ubiquity of these systems, which are found in synchronization processes (e.g., fireflies, power grids, pacemaker cells, coupled oscillators), bird flocking, social epistemology, and opinion dynamics [9, 14, 16, 21, 23, 25, 63]. Agents interact across a time-varying network by averaging their state variables with those of their neighbors. Under mild conditions, such systems are known to converge to a fixed-point attractor.
Of course, the convergence rate of averaging dynamics depends on the underlying network sequences. The case of a single, fixed network is dual to a Markov chain and the convergence rate is given by the corresponding mixing time [41]. When networks are allowed to change over time, bounding the convergence rate becomes far more challenging and suitable analytical tools have been lacking. The introduction of the $s$-energy has given us access to the convergence rates of important instances of averaging dynamics [14]. In this work, we prove new bounds on the $s$-energy, which we then use to bound the convergence rates of systems for bird flocking, opinion dynamics, and distributed motion coordination. In particular, we prove that a group of birds with at most a constant number of flocks converges in polynonomial time; this is the first polynomial bound of its kind. We also provide a theoretical validation for the popular idea of the “Overton Window,” which we interpret as an attracting manifold of viable group opinions.
@@
Our investigation highlights the key role played by network connectivity in the $s$-energy. It helps us resolve what has been a puzzling mystery for several years now: Previous work, indeed, has shown that convergence to equilibrium within $\varepsilon$ is reached in time at most $C\big{(}\log\frac{1}{\varepsilon}\big{)}^{n-1}$, where $n$ is the number of agents and $C$ is a parameter depending only on the networks’ characteristics [16, 29, 37, 55, 56]. When all of the communication networks are connected, however, the time bound drops to $C\log\frac{1}{\varepsilon}$. Such an exponential gap also appears in recent works on broadcast and consensus dynamics and hyper-torpid mixing [17, 20, 26, 27, 65]. We explain this gap by proving that the convergence time is actually of the form $C\big{(}\log\frac{1}{\varepsilon}\big{)}^{m}$, where $m<n$ is the maximum number of connected components at any time. This result follows from our new bounds on the $s$-energy.
@@
  • ran:  a random averaging system is a rev where $q=d\mathbf{1}$ with each $P_{t}$ is the simple random walk over a random $(d-1)$-regular graph $G_{t}$.
@@
The bounds in (a) apply to any initial position $x(0)\in[0,1]^{n}$ or, equivalently, an initial diameter at most 1. For notational convenience, we use a slightly different assumption for the cases (b, c, d): we assume that the initial (scaled) variance $\sum_{i}q_{i}(x_{i}(0)-\mu)^{2}$ is at most 1, where $\mu=\sum_{i}\bar{q}_{i}x_{i}(0)$ is the mean position at time 00; here, $\bar{q}=q/\|q\|_{1}$ is the stationary distribution of $P_{t}$. To see why this different assumption can be made without loss of generality, we note that unit diameter implies a variance of at most $n/\rho$ and, conversely, unit variance implies a diameter bounded by 2.
@@
.absent\!\!\!.\,\,. For constant $c>0$ and any positive $\varepsilon$ small enough,
$\begin{cases}\ \ T_{m,\varepsilon}\leq c(1/\rho)^{n-1}\big{(}mn\log\frac{1}{\varepsilon}\big{)}^{m}\vspace{0.1cm}\\ \ \ T^{\text{\sc rev}}_{m,\varepsilon}\leq c\big{(}\frac{n^{2}}{\rho}\log\frac{1}{\varepsilon}\big{)}^{m}\vspace{0.1cm}\\ \ \ T^{\text{\sc exp}}_{m,\varepsilon}\leq c\big{(}\frac{d^{3}mn}{h^{2}}\log\frac{1}{\varepsilon}\big{)}^{m}\ \ and\ \ \ T^{\text{\sc exp}}_{1,\varepsilon}\leq\frac{cd^{3}}{h^{2}}\log\frac{1}{\varepsilon}\vspace{0.1cm}\\ \ \ \mathbb{E}\,T^{\text{\sc ran}}_{1,\varepsilon}\leq c\log\frac{1}{\varepsilon}\end{cases}$
@@
This paper is organized as follows: Section 2 discusses the applications of Theorem 1.2: bird flocking in §2.1; motion coordination in §2.2; and the Overton window in §2.3. Section 3 proves our $s$-energy bounds: Theorem 1.1 (a) in §3.1; Theorem 1.1 (b) in §3.2; Theorem 1.1 (c) in §3.3; and Theorem 1.1 (d) in §3.4.
2 Applications
2.1 Bird Flocking
Theorem 2.1
.absent\!\!\!.\,\,. A group of $n$ birds forming a maximum of $m\leq n$ flocks relax to within $\varepsilon$ of a fixed velocity vector in time $O(n^{2}/\rho)\log(1/\varepsilon)+t_{o}$, where $t_{o}$ is on the order of $mn^{2(m+2)}(1/\rho)^{m+1}\log\frac{n}{\rho}\,$.
Figure 2: A bird is influenced by its neighbors within distance $r$.
Note that $\bar{q}:=\|q\|_{1}^{-1}q$ is the joint stationary distribution and $q=M_{t}\mathbf{1}\leq\mathbf{1}/\rho$, where $\rho:=\min_{i}a_{i}\in(0,1/2]$. Each nonzero entry in row $i$ of $P_{t}$, being either $a_{i}$ or $1-a_{i}|\mathcal{N}_{i}(t)|\geq a_{i}$, is at least $\rho$. This shows that each one of the three coordinates provides its own reversible averaging system $\mathcal{S}_{j}$ ($j=1,2,3$). The only difference between the systems is their initial states. The $s$-energy of any such system is equal to $\sum_{t}E_{s,t}$, where $E_{s,t}=\sum_{i}l_{i}(t)^{s}$ and $l_{i}(t)$ is the length of the $i$-th block at time $t$. (Recall that a block is an interval formed by the embedded edges.) Let $m$ be the maximum number of flocks and $N_{m,\alpha}$ the number of times $t$ at which some block length $l_{i}(t)$ from at least one of $\mathcal{S}_{j}$ ($j=1,2,3$) is at least $\alpha$. For $10<\alpha<$ 1, we have $N_{m,\alpha}\leq\inf_{s\in(0,1]}3\alpha^{-s}\mathcal{E}_{m,s}^{\text{\sc rev}}$. Our assumption that $\|v_{i}(0)\|\leq\frac{1}{2}\sqrt{\rho/n}$ for all birds $i$ implies that each of the three systems has variance at most 1. Setting $s=c/\log(1/\alpha)$ and assuming that $\alpha$ is small enough so that $0<s\leq 1$, it follows from Theorem 1.1 (b) that, for (another) constant $c>0$,
$N_{m,\alpha}\leq c\Big{(}\frac{n^{2}}{\rho}\log\frac{1}{\alpha}\Big{)}^{m}.$ (2)
2.1.1 Single-flock dynamics
$t_{a}:=(n^{2}/a\rho)^{2},$ (3)
we have
2.1.2 Flock fusion
Lemma 2.2
.absent\!\!\!.\,\,. There is a constant $c=c(\varepsilon_{o})$ such that, for any $i$ and $t>1$,
[…]
  • [1]
    $\bar{w}(t)\leftarrow w_{1}(t)$   and   $l\leftarrow 1$
  • [2]
    for   $k=t-1,\ldots,1$
  • [3]
    if  $\mu_{l}(k)>\alpha$  then  $l\leftarrow\text{argmin}\left\{\,w_{j}(k)\,|\,j\in N_{l}(k)\,\right\}$
  • [4]
    $\bar{w}(k)\leftarrow w_{l}(k)$
[…]
Let $l^{\prime}$ be the value of $l$ in the final assignment $\bar{w}_{l}(1)\leftarrow w_{l}(1)$ in step [4]. Since $\bar{w}(k+1)\geq\bar{w}(k)$ for $k\in R$ and $\bar{w}(t)=w_{1}(t)$, it follows from (6, 8) and $r\leq 1$ that
We set $\alpha=\varepsilon_{o}/t$. Noting that $y_{l^{\prime}}(0)-y_{1}(0)\leq 2$, the lemma follows from (2, 5, 7) and
for constant $d>0$. $\Box$
2.1.3 Stabilization
$t_{o}:=Cmn^{2(m+2)}\left(\frac{1}{\rho}\right)^{m+1}\log\frac{n}{\rho},$
2.2 Distributed Motion Coordination
Let $G$ be a connected (undirected) graph with $n$ vertices labeled in $[n]$ and no self-loops, and let the communication weights $a_{1},\ldots,a_{n}$ be $n$ positive reals such that $a_{i}<1/(d_{i}+1)$, where $d_{i}$ is the degree of vertex $i$. We define $d=\max d_{i}$ and $\rho=\min a_{i}$. For any $t\geq 0$, we define $G_{t}$ by deleting each edge of $G$ with probability $1-p$. We define a (random) stochastic matrix $P_{t}$ for $G_{t}$ as follows:
  1. 1.
    Initialize $P_{t}=0$;
  2. 2.
    If $(i,j)$ is an edge of $G_{t}$, we set $(P_{t})_{ij}=a_{i}$ and $(P_{t})_{ji}=a_{j}$.
  3. 3.
    $(P_{t})_{ii}=1-\sum_{j(j\neq i)}(P_{t})_{ij}$, for all $i$.
Note that every positive entry of $P_{t}$ is at least $\rho$. We embed $G$ in $\mathbb{R}^{3}$ and pin a subset $R$ of $r$ vertices to a fixed plane. We fix the scale by assuming that the embedding lies in the unit cube $[0,1]^{3}$. Without loss of generality, we choose the plane $X=0$. To ensure the immobility of the $r$ vertices, we use symmetrization [14]. This entails attaching to $R$ a copy of $G$ and initializing the embedding of the two copies as mirror-image reflections about $X=0$. This modification increases the number of vertices to $\nu=2n-r$. The sequence $(P_{t})_{t\geq 0}$ is defined by picking a random $G_{t}$ (as defined above) at each step iid.
$D_{t}=\sum_{i}\max_{j:\,(i,j)\in G_{t}}\bigl{(}x_{i}(t)-x_{j}(t)\bigr{)}^{2}.$
By Markov’s inequality,
(a) Initially, moving robots are placed randomly in the unit cube.
(b) Positions of moving robots after 200 steps.
Let $l_{1}(t),\ldots,l_{k_{t}}(t)$ be the lengths of the blocks formed by the edges of $G_{t}$ embedded along the $X$-axis. [^4] Recall that the blocks are the intervals formed by the union of the embedded edges of $G_{t}$. In a slight variant, we define the $s$-energy $E_{s}=\sum_{t\geq 0}E_{s,t}$, where $E_{s,t}=\max_{i=1}^{k_{t}}l_{i}(t)^{s}$. We denote by $W_{s}$ the maximum expected $s$-energy, where the maximum is taken over all initial positions such that $\|x\|_{q}^{2}\leq\nu/\rho$. Since the vertices are embedded in $[-1,1]$ with symmetry about the origin, this applies to the case at hand. Since $q_{i}=1/a_{i}>1$, the initial diameter for $W_{s}$ is at most $\sqrt{2}\,\|x\|_{q}\leq\sqrt{2\nu/\rho}$. (Note that we cannot use 2 as the bound for the diameter since $W_{s}$ only bounds the variance and not the diameter.) By scaling invariance, we have the following recurrence relation:
Let $N_{\alpha}$ be the number of times $t$ at which some block length $l_{i}(t)$ is at least $\alpha$. For $10<\alpha<$ 1, we have $\mathbb{E}\,N_{\alpha}\leq\inf_{s\in(0,1]}\alpha^{-s}W_{s}$. Setting $s=1/\log(n/\rho\alpha^{2})$ yields
Theorem 2.3
.absent\!\!\!.\,\,. The robots align themselves within distance $\epsilon<1$ of a fixed plane in expected time $O\bigl{(}d^{2}n^{4}/p^{3}\rho^{2}\bigr{)}\log(n/\rho\varepsilon)$, where $d$ is the maximum degree of the underlying communication network, $n$ is the number of robots, $1-p$ is the probability of edge failure, and $\rho$ is the smallest communication weight.
2.3 The Overton Window Attractor
  1. 1.
    For $i>n-k$, $(P_{t})_{ii}=1$.
  2. 2.
Theorem 2.4
.absent\!\!\!.\,\,. For any $\delta,\varepsilon>0$, with probability at least $1-\delta$, all of the agents fall within distance $\varepsilon$ of the convex hull of the fixed agents after a number of steps at most
for constant $c>0$.
[…]
$A_{t}=\begin{pmatrix}Q_{t}&v\\ \mathbf{0}_{h}^{T}&1\end{pmatrix}.$
We define the row vector $u\in\mathbb{R}^{h}$ by setting its $i$-th coordinate to $\rho$ if $v_{i}>0$ and 00 otherwise. We require that $1-2\|u\|_{1}\geq\rho$; hence $\rho\leq 1/(2d_{t}+1)$, where $d_{t}$ is the number of mobile agents (among the $h$ of them) adjacent in $G_{t}$ to at least one fixed agent. This condition is easily satisfied by setting $\rho\leq 1/2n$. The evolution follows the update: $z(t+1)=B_{t}\,z(t)$, where $z(t)\in[-1,1]^{\nu}$ and $z_{h+1}(0)=0$.
Let $G^{*}$ be the augmented $\nu$-vertex graph formed from $G$ and let $G_{t}^{*}$ be its subgraph selected at time $t$. Note that, via $z(t)$, these graphs are embedded in $[-1,1]^{\nu}$. If $\Delta_{t}$ denotes the length of the longest edge of $G^{*}$ at time $t$ and $T_{\alpha}$ is the last time at which the diameter of the system is at least $\alpha$, then $\Delta_{t}\geq\alpha/\nu$ for all $t\leq T_{\alpha}$ because $G^{*}$ is connected. The longest edge in $G^{*}$ (with ties broken alphabetically) appears in $G_{t}^{*}$ with probability at least $p$. Fix $s\in(0,1]$ and define the random variable $\chi_{t}$ to be $\Delta_{t}^{s}$ if the longest edge of $G$ at time $t$ is in $G_{t}$ and 00 otherwise. By [16], the maximum $s$-energy is at most $\mathcal{E}_{s}\leq 2^{s}(3/\rho s)^{\nu-1}$; hence
Minimizing the right-hand side over all $s\in(0,1]$ yields
[…]
$q:=\sum_{j=1}^{h}(P_{\leq t})_{ij}=\sum_{j=1}^{h}(Q_{\leq t})_{ij}\leq\alpha.$
Trivially, $x_{i}(t+1)=q\tilde{u}+(1-q)\tilde{v}$, where
Observing that $\tilde{v}$ lies in the convex hull of the fixed agents, we form the difference $x_{i}(t+1)-\tilde{v}=q(\tilde{u}-\tilde{v})$ and note that the distance from $x_{i}(t+1)$ to the hull is bounded by $q\|\tilde{u}-\tilde{v}\|_{2}\leq\alpha\sqrt{d}$. Setting $\alpha=\varepsilon/\sqrt{d}$ completes the proof. $\Box$
Theorem 2.5
.absent\!\!\!.\,\,. For any $\delta,\varepsilon>0$, with probability at least $1-\delta$, all of the agents fall within distance $\varepsilon$ of the convex hull of the fixed agents in time bounded by
for constant $c>0$, from which the theorem follows. $\Box$
3 The Proofs
@@
for any $i$ in $[u,v]$ and $y_{i}=x_{i}$ otherwise. We repeat this step indefinitely. Twist systems are highly nondeterministic. At each step, a new interval $[u,v]\subseteq[n]$, called a block, is picked and the agents’ motion is only constrained by (11) and the need to maintain their ranks (ie, agents never cross). It is not entirely obvious that twist systems should always exist: in other words, that the constraints imposed in (11) are always satisfiable. This is a direct consequence of Lemma 3.1, which we state and prove below.
For the purposes of this work, we extend the concept to $m$-twist systems by stipulating, at each time $t$, a partition of $[n]$ into up to $m_{t}$ blocks $[u_{t,l},v_{t,l}]$ ($1\leq l\leq m_{t}\leq m$). Each agent is now subject to (11) within its own enclosing block. We define the $s$-energy $E_{s}^{\text{\sc tw}}$ of a twist system as we did with an averaging system by adding together the $s$-th powers of all the block lengths. We use the same notation with the addition of the superscript tw.
@@
.absent\!\!\!.\,\,. An averaging system with at most $m$ connected components at any time can be interpreted as an $m$-twist system with the same $s$-energy.
Proof. Fix an averaging system and let $(x_{i})_{i=1}^{n}$ and $(y_{i})_{i=1}^{n}$ be the positions of the agents at times $t$ and $t+1$, given in nondecreasing order. We denote by $x_{i}^{\prime}$ the position of agent $i$ at time $t+1$. Let $[x_{u},x_{v}]$ be a block at time $t$ (ie, an interval of the union of the embedded edges of $G_{t}$). Pick $k<v$ and write $z=\rho x_{k}+(1-\rho)x_{v}$. All the diagonal elements of $P_{t}$ are at least $\rho$; hence $x^{\prime}_{i}\leq z$, for all $i\leq k$, and $y_{k}\leq z$. In fact, we even have $y_{k+1}\leq z$. Indeed, the embedded edges of $G_{t}$ cover all of $[x_{u},x_{v}]$, so at least one of them, call it $(l,r)$, must join $[u,k]$ to $[k+1,v]$; hence $x^{\prime}_{r}\leq z$. Our claim follows. This proves that, for all $i\in(u,v]$, $y_{i}\leq\rho x_{\max\{i-1,u\}}+(1-\rho)x_{v}$. We omit the case $i=u$ and the mirror-image inequality, which repeat the same argument. Summing up all the powers $(x_{v}-x_{u})^{s}$ shows the equivalence between the two $s$-energies. $\Box$
@@
.absent\!\!\!.\,\,. For any $z\geq 1$, $\mathcal{E}_{m,1}^{\text{\sc tw}}\leq 2z^{-\nu}\,\mathcal{K}(z)$, where $\nu=\lceil n/2\rceil$.
@@
We consider the case $m>1$. Fix $t\leq T_{\delta}$; if $m_{t}>1$, let $j$ maximize $x_{j+1}(t)-x_{j}(t)$ over all $j=v_{i}$ and $i<m_{t}$ (break ties by taking the smallest $j$). This corresponds to the maximum distance between consecutive blocks. For this reason, we call $(j,j+1)$ the max-gap at time $t$. We say that $t$ is ungapped if $m_{t}=1$ or $x_{j+1}(t)-x_{j}(t)\leq\delta/m$; it is gapped otherwise. Assuming that $t$ is gapped, let $(j,j+1)$ be its max-gap and write $\zeta_{t}=\min_{k}\bigl{\{}\,t<k\leq T_{\delta}\,|\,\exists\,l:u_{k,l}\leq j<v_{k,l}\,\bigr{\}}$: If the set is empty, we set $\zeta_{t}=T_{\delta}$; else $l$ is unique and we denote it by $l_{t}$. We call the interval $[t,\zeta_{t}]$ a span and the block $l_{t}$, if it exists, its cap. We note that $x_{j+1}(k)-x_{j}(k)$ cannot decrease during the times $k=t,\ldots,\zeta_{t}$. This shows that a cap covers a length greater than $\delta/m$.
We begin with a few words of intuition. The energetic contribution of an ungapped time $t$ is easy to account for: It is at most $m$. On the other hand, the $1$-energy is at least the diameter minus the added length of the gaps between blocks, which amounts to at least $31-2\delta\geq 1/$ 3; in other words, $E_{s,t}\leq 3mE_{1,t}$. Summing up over all ungapped times and plugging in our bound for $s=1$ gives us the desired result. Accounting for gapped times is more difficult, as it requires dealing with small scales. If we had only one span, we could simply split the system into two decoupled subsystems and set up a recurrence relation. The problem is that the presence of $k$ capped spans would force us to repeat the recursion $k-1$ times. With no apriori bound on $k$, this approach is not too promising. Instead, we make a bold move: We argue that, because a cap is longer than $\delta/m$, its own 1-energy contribution (ie, its length) is large enough to “pay” for the $s$-energy of its entire span. This is not quite right, of course, but one can fix the argument by using the weighted 1-energy of the cap and upscaling it suitably. Once again, this reduces the problem to the case $s=1$, so our method is, in effect, a linearization. Here is the proof.
@@
[…]
  • 1.
    $\mathcal{G}\leftarrow\emptyset$ and $t^{\prime}\leftarrow 0$;
  • 2.
    if $t\leftarrow\min\,\bigl{\{}\,\text{gapped }i\,|\,t^{\prime}\leq i\leq T_{\delta}\,\bigr{\}}$ exists
  • 3.
    then $\mathcal{G}\leftarrow\mathcal{G}\cup\bigl{\{}\,i\,|\,t\leq i\leq\zeta_{t}\,\bigr{\}}$;
  • 4.
    if $\zeta_{t}<T_{\delta}$ then $t^{\prime}=\zeta_{t}+1$; go to 2;

### v2

@@ Abstract
Many multi-agent systems evolve by repeatedly updating each state to a weighted average of its neighbors, a process known as averaging dynamics, whose behavior becomes difficult to analyze when the interaction network varies over time. In recent years, the $s$-energy has emerged as a useful tool for bounding the convergence rates of such systems, complementing the classical techniques that rely on fixed graphs. We derive new bounds on the $s$-energy under minimal connectivity assumptions. As a consequence, we obtain convergence guarantees for several models of collective dynamics and resolve a number of open questions in the areas. Our results highlight the dependence of the $s$-energy on the connectivity of the underlying networks and use it to explain the exponential gap in the convergence rates of stationary and time-varying consensus systems.
@@ 1 Introduction
We study discrete-time averaging systems defined by sequences of stochastic matrices whose nonzero entries specify an underlying, possibly time-varying interaction graph. At each step, agents update their states by averaging them with those of their neighbors. Such systems arise naturally when local interactions are encoded by networks and global evolution is governed by repeated application of averaging operators [7, 9, 10, 11, 12, 13, 29, 34, 35, 36, 37, 39, 40, 43, 44, 45, 46, 49, 51, 52, 55, 62, 14, 16, 21, 23, 25, 63]. From a mathematical perspective, these dynamics interpolate between products of stochastic matrices, random walks on graphs, and time-inhomogeneous Markov chains.
When the interaction graph is fixed, the evolution reduces to a random walk governed by a single stochastic matrix. Its convergence properties are well understood through spectral graph theory and mixing time [41]. By contrast, once the graph and weights are allowed to change over time, classical tools based on eigenvalues or stationary distributions no longer work. Understanding convergence in this setting therefore requires new invariants that remain meaningful under arbitrary graph sequences subject only to local constraints.
One such invariant is the $s$-energy, a scale-sensitive function associated with the evolution of averaging systems [14]. Defined independently of any particular application, the $s$-energy measures how pairwise distances between agents shrink over time across multiple spatial scales. The parameter $s\in(0,1]$ controls the scale at which distances are emphasized, enabling the function to detect both global and local behavior.
Our main results establish new bounds on the $s$-energy under minimal assumptions on the interaction weights and graph structure. In particular, for systems whose interaction graph has at most $m$ connected components at every time step, we show that the total $s$-energy is bounded by a function that is polynomial in the number of agents and exponential only in $m$. The bounds hold for arbitrary time-varying graphs, and improve further under additional assumptions.
@@
Our result highlights the key role played by network connectivity in the $s$-energy. It helps us resolve what has been a puzzling mystery for several years now: Previous work, indeed, has shown that convergence to equilibrium within $\varepsilon$ is reached in time at most $C\big(\log\frac{1}{\varepsilon}\big)^{n-1}$, where $n$ is the number of agents and $C$ is a parameter depending only on the networks’ characteristics [16, 29, 37, 55, 56]. When all of the communication networks are connected, however, the time bound drops to $C\log\frac{1}{\varepsilon}$. Such an exponential gap also appears in recent works on broadcast and consensus dynamics and hyper-torpid mixing [17, 20, 26, 27, 65]. We explain this gap by proving that the convergence time is actually of the form $C\big(\log\frac{1}{\varepsilon}\big)^{m}$, where $m<n$ is the maximum number of connected components at any time. This result follows from our new bounds on the $s$-energy.
The abstract nature of the $s$-energy framework allows it to be applied uniformly across a range of models. As corollaries to our main bounds, we obtain polynomial-time convergence guarantees for multi-flock bird dynamics, attraction rates for opinion dynamics with stubborn agents, an explanation of the “Overton window” phenomenon, as well as bounds for several classes of randomized or reversible averaging processes. In each case, the system-specific analysis reduces to verifying that the dynamics fit within the general averaging framework.
From a broader perspective, the $s$-energy can be viewed as a nonlinear analogue of classical energies used in the analysis of random walks and graph Laplacians. While spectral graph and Dirichlet forms quantify contraction in fixed graphs, the $s$-energy remains well behaved under graph changes, adversarial updates, and non-reversible dynamics. To our knowledge, it provides the first general-purpose tool for bounding convergence in time-varying averaging systems without imposing global regularity conditions.
The paper is organized as follows. Section 2 introduces the averaging framework, formally defines the $s$-energy, and states our results mathematically. Section 3 proves main results: general upper and lower bounds. Section 4 applies the theory to several dynamical systems arising in distributed control, opinion dynamics, and collective motion.
2 Preliminaries and Models
We now formalize the class of averaging systems studied in this paper and introduce the notation and assumptions used throughout. An averaging system consists of a sequence of stochastic matrices $(P_{t})_{t\geq 0}$, where each $P_{t}$ specifies a graph $G_{t}$ in the sense that $(P_{t})_{ij}>0$ if and only if $(i,j)$ is an edge of $G_{t}$. The state vector $x(t)\in\mathbb{R}^{n}$ then evolves according to
$x(t+1)=P_{t}x(t).$
Within this framework, we define the $s$-energy, a scale-sensitive functional associated with the evolution of the system, and recall several basic conventions that will be used throughout.
@@ 2 Preliminaries and Models
  • ran: a random averaging system is a rev where $q=d\mathbf{1}$ with each $P_{t}$ is the simple random walk over a random (d1)(d-1)-regular graph $G_{t}$.
@@
The bounds in (a) apply to any initial position $x(0)\in[0,1]^{n}$ or, equivalently, an initial diameter at most 1. For notational convenience, we use a slightly different assumption for the cases (b, c, d): we assume that the initial (scaled) variance $\sum_{i}q_{i}(x_{i}(0)-\mu)^{2}$ is at most 1, where $\mu=\sum_{i}\bar{q}_{i}x_{i}(0)$ is the mean position at time 0; here, $\bar{q}=q/\|q\|_{1}$ is the stationary distribution of $P_{t}$. To see why this different assumption can be made without loss of generality, we note that unit diameter implies a variance of at most $n/\rho$ and, conversely, unit variance implies a diameter bounded by 2.
@@
.\!\!\!.\,\, For constant $c>0$ and any positive $\varepsilon$ small enough,
$\begin{cases}\ \ T_{m,\varepsilon}\leq c(1/\rho)^{n-1}\big(mn\log\frac{1}{\varepsilon}\big)^{m}\vskip 2.84544pt\\ \ \ T^{\text{\sc rev}}_{m,\varepsilon}\leq c\big(\frac{n^{2}}{\rho}\log\frac{1}{\varepsilon}\big)^{m}\vskip 2.84544pt\\ \ \ T^{\text{\sc exp}}_{m,\varepsilon}\leq c\big(\frac{d^{3}mn}{h^{2}}\log\frac{1}{\varepsilon}\big)^{m}\ \ and\ \ \ T^{\text{\sc exp}}_{1,\varepsilon}\leq\frac{cd^{3}}{h^{2}}\log\frac{1}{\varepsilon}\vskip 2.84544pt\\ \ \ \mathbb{E}\,T^{\text{\sc ran}}_{1,\varepsilon}\leq c\log\frac{1}{\varepsilon}\end{cases}$
@@
3 Energy Bounds for Averaging Systems
In this section, we prove Theorem 2.1. The four parts correspond to different structural assumptions on the averaging system. Part (a) is the most general case and contains the core argument of the paper. Parts (b)-(d) introduce additional structure that allows alternative techniques and leads to sharper bounds.
@@ 3 Energy Bounds for Averaging Systems
for any $i$ in $[u,v]$ and $y_{i}=x_{i}$ otherwise. We repeat this step indefinitely. Twist systems are highly nondeterministic. At each step, a new interval $u,v]\subseteq[n]$, called a block, is picked and the agents’ motion is only constrained by (1) and the need to maintain their ranks (ie, agents never cross). It is not entirely obvious that twist systems should always exist: in other words, that the constraints imposed in (1) are always satisfiable. This is a direct consequence of Lemma 3.1, which we state and prove below.
For the purposes of this work, we extend the concept to $m$-twist systems by stipulating, at each time $t$, a partition of $[n]$ into up to $m_{t}$ blocks $u_{t,l},v_{t,l}]$ ($1\leq l\leq m_{t}\leq m$). Each agent is now subject to (1) within its own enclosing block. We define the $s$-energy $E_{s}^{\text{\sc tw}}$ of a twist system as we did with an averaging system by adding together the $s$-th powers of all the block lengths. We use the same notation with the addition of the superscript tw.
@@
.\!\!\!.\,\, An averaging system with at most $m$ connected components at any time can be interpreted as an $m$-twist system with the same $s$-energy.
Proof. Fix an averaging system and let $(x_{i})_{i=1}^{n}$ and $(y_{i})_{i=1}^{n}$ be the positions of the agents at times $t$ and $t+1$, given in nondecreasing order. We denote by $x_{i}^{\prime}$ the position of agent $i$ at time $t+1$. Let $x_{u},x_{v}]$ be a block at time $t$ (ie, an interval of the union of the embedded edges of $G_{t}$). Pick $k<v$ and write $z=\rho x_{k}+(1-\rho)x_{v}$. All the diagonal elements of $P_{t}$ are at least $\rho$; hence $x^{\prime}_{i}\leq z$, for all $i\leq k$, and $y_{k}\leq z$. In fact, we even have $y_{k+1}\leq z$. Indeed, the embedded edges of $G_{t}$ cover all of $x_{u},x_{v}]$, so at least one of them, call it $(l,r)$, must join $[u,k]$ to $[k+1,v]$; hence $x^{\prime}_{r}\leq z$. Our claim follows. This proves that, for all $i\in(u,v]$, $y_{i}\leq\rho x_{\max\{i-1,u\}}+(1-\rho)x_{v}$. We omit the case $i=u$ and the mirror-image inequality, which repeat the same argument. Summing up all the powers $(x_{v}-x_{u})^{s}$ shows the equivalence between the two $s$-energies. $\Box$
@@
.\!\!\!.\,\, For any $z\geq 1$, $\mathcal{E}_{m,1}^{\text{\sc tw}}\leq 2z^{-\nu}\,\mathcal{K}(z)$, where $\nu=\lceil n/2\rceil$.
@@
We consider the case $m>1$. Fix $t\leq T_{\delta}$; if $m_{t}>1$, let $j$ maximize $x_{j+1}(t)-x_{j}(t)$ over all $j=v_{i}$ and $i<m_{t}$ (break ties by taking the smallest $j$). This corresponds to the maximum distance between consecutive blocks. For this reason, we call $(j,j+1)$ the max-gap at time $t$. We say that $t$ is ungapped if $m_{t}=1$ or $x_{j+1}(t)-x_{j}(t)\leq\delta/m$; it is gapped otherwise. Assuming that $t$ is gapped, let $(j,j+1)$ be its max-gap and write $\zeta_{t}=\min_{k}\bigl\{\,t<k\leq T_{\delta}\,|\,\exists\,l:u_{k,l}\leq j<v_{k,l}\,\bigr\}$: If the set is empty, we set $\zeta_{t}=T_{\delta}$; else $l$ is unique and we denote it by $l_{t}$. We call the interval $t,\zeta_{t}]$ a span and the block $l_{t}$, if it exists, its cap. We note that $x_{j+1}(k)-x_{j}(k)$ cannot decrease during the times $k=t,\ldots,\zeta_{t}$. This shows that a cap covers a length greater than $\delta/m$.
We begin with a few words of intuition. The energetic contribution of an ungapped time $t$ is easy to account for: It is at most $m$. On the other hand, the $1$-energy is at least the diameter minus the added length of the gaps between blocks, which amounts to at least $1-2\delta\geq 1/3$; in other words, $E_{s,t}\leq 3mE_{1,t}$. Summing up over all ungapped times and plugging in our bound for $s=1$ gives us the desired result. Accounting for gapped times is more difficult, as it requires dealing with small scales. If we had only one span, we could simply split the system into two decoupled subsystems and set up a recurrence relation. The problem is that the presence of $k$ capped spans would force us to repeat the recursion k1k-1 times. With no apriori bound on $k$, this approach is not too promising. Instead, we make a bold move: We argue that, because a cap is longer than $\delta/m$, its own 1-energy contribution (ie, its length) is large enough to “pay” for the $s$-energy of its entire span. This is not quite right, of course, but one can fix the argument by using the weighted 1-energy of the cap and upscaling it suitably. Once again, this reduces the problem to the case $s=1$, so our method is, in effect, a linearization. Here is the proof.
@@
       1. $\mathcal{G}\leftarrow\emptyset$ and $t^{\prime}\leftarrow 0$; 2. if $t\leftarrow\min\,\bigl\{\,\text{gapped }i\,|\,t^{\prime}\leq i\leq T_{\delta}\,\bigr\}$ exists 3. then $\mathcal{G}\leftarrow\mathcal{G}\cup\bigl\{\,i\,|\,t\leq i\leq\zeta_{t}\,\bigr\}$; 4. if $\zeta_{t}<T_{\delta}$ then $t^{\prime}=\zeta_{t}+1$; go to 2;